# Perry: A High-level Framework for Accelerating Cyber Deception Experimentation

### v1

@@ Abstract
Cyber deception aims to distract, delay, and detect network attackers with fake assets such as honeypots, decoy credentials, or decoy files. However, today, it is difficult for operators to experiment, explore, and evaluate deception approaches. Existing tools and platforms have non-portable and complex implementations that are difficult to modify and extend. We address this pain point by introducing Perry, a high-level framework that accelerates the design and exploration of deception what-if scenarios. Perry has two components: a high-level abstraction layer for security operators to specify attackers and deception strategies, and an experimentation module to run these attackers and defenders in realistic emulated networks. To translate these high-level specifications we design four key modules for Perry: 1) an action planner that translates high-level actions into low-level implementations, 2) an observability module to translate low-level telemetry into high-level observations, 3) an environment state service that enables environment agnostic strategies, and 4) an attack graph service to reason about how attackers could explore an environment. We illustrate that Perry’s abstractions reduce the implementation effort of exploring a wide variety of deception defenses, attackers, and environments. We demonstrate the value of Perry by emulating 55 unique deception what-if scenarios and illustrate how these experiments enable operators to shed light on subtle tradeoffs.
@@ I Introduction
Cyber deception defenses seek to deceive an attacker with fake resources, information, or critical assets. These could include deploying a decoy host (e.g., a honeypot) [47, 45, 40], decoy credentials [60, 20], decoy files [20], or other system assets. Cyber deception holds promise to distract attackers [16, 20], detect attacks in progress [60, 22], and delay attackers from executing attacks [16]. With emerging AI-based attackers [11, 51], there is speculation that cyber deception may become a critical part of the defense arsenal [29].
In order to inform their future security posture, operators would like the ability to run a broad spectrum of deception what-if scenarios relevant to their environment. For example: Would deploying many decoy files to a network distract attackers? Where are the optimal locations to deploy decoy credentials in a network? From these experiments, security operators could collect quantitative measurements on cost (e.g., number of decoys used) and effectiveness (e.g., data exfiltrated, hosts infected).
Unfortunately, running such experiments with existing tools such as Caldera and Elastic Security [5, 30, 56, 23, 25, 13] entails significant effort and complexity. We find that current tools for specifying attacker (e.g., execute exploits, command infected computers) and deception (e.g., react to fake credentials) strategies operate at a very low level of abstraction. These tools rely on low-level mechanisms such as command-line interfaces, manual configuration files, and raw network log parsing, making them complex and cumbersome to use. The lack of abstraction and modularity also means that attacker and defender implementations entail complex and distributed changes. Thus, operators cannot easily take existing implementations of strategies and easily port it to work in their own environments and experiments. Taken together, this means that operators are stymied in their efforts to run deception-related experiments.
@@
To address this challenge, we present Perry, a high-level framework for accelerating cyber deception experimentation. Perry has two components: (1) a high-level abstraction layer for security operators to specify attackers and deception defenders in a portable environment-agnostic manner, and (2) an experimentation module to run these attackers and defenders in realistic emulated networks (Fig. 1). Operators use a high-level and environment-agnostic state machine abstraction to specify attack and deception strategies, that execute high-level actions based on high-level observations of the environment (Sec. III). Because these specifications are environment agnostic, they are naturally portable from one setting to another; e.g., operators can take open source contributions of existing attacker and defender strategies (both from Perry developers and the broader community) and run them seamlessly in their own emulated environments (with comparable telemetry and actuation capabilities) with zero effort.
Perry uses four key modules to translate high-level attacker and deception strategies into low-level primitives: 1) an action translator to specify attack and defense actions at a high level (e.g., deploy a decoy), 2) an event-driven observability module that emits high-level observations (e.g., host is suspicious), 3) an environment state service to enable environment agnostic strategies, and 4) an attack graph service to reason about how an attacker can explore an environment.
We show that Perry’s abstractions can reduce the implementation effort across a wide variety of defense and attack strategies (see Sec. VII-A). We validate that there is no effort needed to port attacker and deception strategies across environments due to the environment-agnostic and portable specifications. We quantify the reduction in lines of code (LOC) by 17–27$\times$ for attacker implementations and 7–14$\times$ for deception implementations, and 3) optionally using LLM code generation to create correct implementations of deception strategies in Perry. For instance, in Sec. VII-A we illustrate evaluating two attackers and two deception approaches across five environments. With Perry, users only need to write each attack and deception strategy once, and these can be applied to all environments without additional effort. In contrast, previous frameworks would require users to manually re-implement each attack and defense strategy for each environment. In Sec. VII-A we also measure the LOC of deception approaches and attackers with and without using Perry’s abstractions.
This reduction in effort can enable operators to rapidly 1) explore a variety of deception strategies against different attackers in different environments (see Sec. VII-A), 2) implement variations of strategies (see Sec. VII-B), and 3) evaluate extensions to consider new defense and attacker capabilities (see Sec. VII-B). Using Perry, we explore diverse and scenarios and can shed light on the efficacy of deceptive defenses (see Sec. VII-C). For instance,
@@
    We introduce a novel high-level and environment agnostic abstraction for attackers and deception defenders: a state machine that executes high-level actions based on high-level observations of the environment (Sec. III).
@@
    We demonstrate the value of Perry by illustrating how Perry can enable operators to rapidly explore a wide variety of deception defenses, attackers, environments, and can shed light on the tradeoffs of deception defenses(Sec. VII).
Reproducible Research Perry is open-source and will be publicly available to the research community.
@@ II-A The case for what-if experiments in deception
Imagine an attacker is exploring a network, if and how, could deception techniques have prevented the data breach: Would some deception strategies have higher efficacy than others? To shed light on these questions, operators may want to explore what-if scenarios:
Example 1. The operator wants to quickly explore the efficacy of a wide variety of deception capabilities (e.g., decoy hosts [45], fake credentials [60], or fake software patches [6]) against against a variety of attackers.
Example 2. The operator may want to understand the benefits of adding better telemetry with the above deception algorithms, enabling the defense to react to attackers actions. For example, an operator may want to detect and react to attackers interacting with their fake credentials.
@@ II-B Requirements and limitations of prior work
    Extensible: As new deception approaches, system capabilities, and attack strategies appear, the platform should be extensible to evaluate these new techniques.
At one extreme, we can consider simple “toy” agent-based simulators that are low effort [39, 28, 10, 57]. These do not emulate real networks, attackers or defenders and lack the realism and expressiveness for real-world scenarios. We will later see in Sec. VII-A, subtle implementation details can greatly impact the efficacy of deception.
@@
Specifically, we consider a stateful defender (Example 2) that 1) deploys decoy hosts and credentials to the network, 2) detects hosts interacting with the decoy hosts and using the decoy credentials, and 3) restores any hosts that interact with decoy resources. Similarly, we create an attacker that mimics the real-world Equifax attacker [36] by: 1) gains initial access to the network by infecting vulnerable web servers, 2) searches and uses credentials to infect database hosts, and 3) exfiltrates data from each database host. Fig. 3 shows code snippets of the defender implementation.
Next, we highlight three key issues with such implementations.  [^1] Our goal is to illustrate problems broadly with available tooling, not at these specific tools. For instance, the problems we highlight with Caldera also exist in other attacker emulation tools [48, 23, 15] and the pain points of Elasticsearch with Python are similar to other defense tools [8, 25, 55].)
@@
    Low-level and cumbersome code: Existing tools offer low-level SDKs resulting in cumbersome low-level code. For instance, the stateful defender implementation requires 272 lines of low-level code. Deploying decoy hosts and credentials requires 213 lines of code using the Ansible and Openstack SDKs. The remaining 59 lines of code use the Elasticsearch API to detect detect attackers interacting with decoy hosts and credentials.
@@
    Non-portable code: Existing tools force the code to be tightly coupled with the environment, making it difficult to port attackers and defenders to the operator’s environment. For instance, the attacker implementation exfiltrates data. In the Equifax environment, this requires (1) choosing a correct path of two hosts to exfiltrate the data on, and (2) choose the correct exfiltration protocol between the hosts (e.g., the web server communicates with the database instance with SSH while the attacker and web server communicate with HTTPS). If the operator’s environment has a different path or different protocols, the attacker implementation will fail.
@@
    Modifications require distributed changes: As a consequence of complex low-level code, existing tools do not enforce modularity of strategies, capabilities, or telemetry making it difficult to extend to evaluate new deception approaches and attackers. Our implementations’ strategies became tightly coupled to their capabilities and telemetry making them difficult to extend. For instance, we may want to extend our deception to add decoy files to decoy hosts. But adding decoy files requires many distributed changes to the implementation. Specifically, In Fig. 3, we have to: (1) extend the initialization function to maintain lists of real and decoy file, (2) extend the detection rules to detect if fake data is being exfiltrated, (3) and (4) implement Ansible Playbooks to deploy the decoy files.
@@ III Perry System Overview
In this section, we give an overview of Perry. Perry introduces a high-level abstraction layer for security operator to create attackers and deception defenders, and an experimentation module to emulate these attackers and defenders in realistic network topologies.
@@
The input to Perry is the environment, attacker, and defender, shown in Fig. 4. Perry first instantiates the environment on a virtualized network, and executes the defender and attacker logic expressed using its high-level abstraction. Perry collects metrics such as defender resources used (e.g., number of decoys used) and attacker success metrics (e.g., data exfiltrated) to enable quantitative analytics (Sec. VII).
@@ III-A High-level idea
To put the key ideas on Perry in context, let us revisit from first principles what the operator was trying to do and why existing tools failed to provide the right capabilities for expressing those intents.
Conceptually, the operator has a mental model of the attacker and defender logic in Sec. II shown in the simple state machines in Fig. 5. Here, the defender wants to deploy decoys, wait for attackers to interact with the decoys, and then runs actions to stop the attacker. Similarly, the attacker runs like a logical state machine that tries to infect web servers, find credentials, and finally exfiltrate data.
While these intents are easy to describe in this mental model, there is a significant disconnect with existing tools. Existing frameworks are environment specific, low-level and lack any abstraction, making it painful to realize these intents. For example, the first state of the defender in Fig. 5 is to deploy decoy hosts and credentials. But this state in our implementation requires 240 lines of code (163 lines of Ansible code and 77 lines of Python code) and only works in one environment. As we discussed earlier, this code is tightly coupled, making it difficult to extend to new environments, capabilities, and strategies.
@@
Based on this insight, Perry introduces a high-level programming model where we explicitly represent attackers and defenders as logical state machines that execute high-level actions based on high-level observations of the environment. In essence, Perry helps the user express high-level intentions of what they need to do and decoupling it from the how to express it in low-level tools.
@@ III-B Anatomy of an experiment
Deception defense: In Perry, deception defense has three stages: A) initialization, strategies setup and deploy deception resources, B) event listeners, strategies subscribe to high-level observations from the telemetry service, and C) event handlers, strategies react to the high-level observations. For instance, the pseudo code of the stateful deception approach in Fig. 5 is (full example in Sec. V):
@@
# (B) Event listners, detect interactions e.g.,
@@ IV Detailed system design
Next, we describe the design of the key modules in Perry: an environment state service, an action translator, an observability module, and an attack graph service. Broadly speaking, strategies subscribe to high-level observations from the observability moduleand output high-level actions to be translated by the action translator, shown in Fig. 6.
@@ IV-A Environment state service
As shown in Sec. II, strategies in existing tools are often specific to a single environment. Rather than strategies having low-level and environment specific code, in Perry, code that reasons about the environment is delegated to the environment state service. Perry modules and strategies query the environment state service for environment information (e.g., check if a credential is fake). To answer these queries, the environment state service maintains a network information base [23] by subscribing to high-level events from the observability module (Sec. IV-C). For example, for attackers, the environment state service subscribes to a new host discovered event and adds hosts to the knowledge base as they are discovered.
As an illustration, the following snippet shows a simple randomized decoy host and credential deployment using the Perry API. The key idea here is that we delegate all environment specific code to the environment state service:
@@
In the above example, the implementation is now environment agnostic. Rather than the code specifying low-level individual IP addresses to deploy decoy credentials, it uses the environment state service to select the high-level intention of the host to deploy the decoy credentials on.
@@ IV-B Action translator
Currently, actions in attack and defense emulation tools [5, 23, 15, 49] are specified with low-level commands that have pre and post-conditions. For example, one command might be an nmap scan script with the post-condition to execute an SSH exploit if a CVE is found.
@@
In Perry, the high-level actions are decoupled from there implementations. As a result, the action translator makes it easier to specify, modify and reason about strategies because users no longer have to reason about low-level code. Furthermore, the decoupling of implementations makes the action translator extensible in terms of both adding new high-level actions (e.g., adding a stealthy scan action) and adding new implementations of actions (e.g., a better data exfiltration technique). For example, in a strategy we can program an exfiltrate data action as:
@@
We also design a library of common high-level deception and attack actions that are portable across environments in Table I. For instance, the data exfiltration action (1) uses the attack graph service to identify the shortest exfiltration path, and (2) uses the environment state service to choose the correct protocol (e.g., HTTP) to use when exfiltrating the data.

### v2

@@ Abstract
Cyber deception aims to distract, delay, and detect network attackers with fake assets such as honeypots, decoy credentials, or decoy files. However, today, it is difficult for operators to experiment, explore, and evaluate deception approaches. Existing tools and platforms have non-portable and complex implementations that are difficult to modify and extend. We address this pain point by introducing Perry, a high-level framework that accelerates the design and exploration of deception what-if scenarios. Perry has two components: a high-level abstraction layer for security operators to specify attackers and deception strategies and an experimentation module to run these attackers and defenders in realistic emulated networks. To translate these high-level specifications into low-level primitives, we design four key modules in Perry: 1) an action planner that translates high-level actions into low-level implementations, 2) an observability module to translate low-level telemetry into high-level observations, 3) an environment state service that enables environment agnostic strategies, and 4) an attack graph service to reason how attackers could explore an environment. We illustrate that Perry’s abstractions reduce the implementation effort across a wide variety of deception defenses, attackers, and environments. We demonstrate the value of Perry by emulating 55 unique deception what-if scenarios, illustrating how these experiments enable operators to shed light on subtle tradeoffs.
@@ I Introduction
Cyber-deception-based defenses seek to deceive an attacker with fake resources, information, or critical assets. These could include a decoy host (e.g., a honeypot) [50, 48, 42], decoy credentials [64, 22], decoy files [22], or other system assets. Cyber deception holds promise to distract attackers [17, 22], detect attacks in progress [64, 24], and delay attackers [17]. With emerging AI-based attackers [12, 54], cyber deception may become a critical part of the defense arsenal [31].
In order to inform their future security posture, security operators need the ability to run a broad spectrum of deception what-if scenarios relevant to their environment. For example: Would deploying many decoy files to a network distract attackers? Where are the optimal locations to deploy decoy credentials in a network? From these experiments, security operators could collect quantitative measurements on cost (e.g., number of decoys used) and effectiveness (e.g., data exfiltrated, hosts infected).
Unfortunately, running such experiments with existing tools such as Caldera and Elastic Security [5, 32, 60, 25, 27, 14] entails significant operator effort and complexity. We find that current tools for specifying attacker (e.g., execute exploits, command infected computers) and deception (e.g., detect fake credential usage) strategies operate at a very low level of abstraction. These tools rely on low-level mechanisms such as command-line interfaces, manual configuration files, and raw network log parsing, making them complex and cumbersome to use. The lack of abstraction and modularity also means that changes to attacker and defender implementations entail complex and distributed changes. Thus, operators cannot easily take existing implementations of strategies and quickly port it to work in their own environments and experiments. Taken together, this means that operators are stymied in their efforts to run deception-related experiments.
@@
To address this challenge, we present Perry, a high-level framework for accelerating cyber deception experimentation. Perry has two components: (1) a high-level abstraction layer for security operators to specify attackers and deception defenders in a portable environment-agnostic manner and (2) an experimentation module to run these attackers and defenders in realistic emulated networks (Fig. 1). Operators use a high-level and environment-agnostic state machine abstraction to specify attack and deception strategies, which execute high-level actions based on high-level observations of the environment (Sec. III). Because these specifications are environment agnostic, they are naturally portable from one setting to another; for example, operators can take open-source contributions of existing attacker and defender strategies (either from Perry developers and the broader community) and run them seamlessly in their own emulated environments (with comparable telemetry and actuation capabilities) with low effort.
Perry uses four key modules to translate high-level attacker and deception strategies into low-level primitives: 1) an action translator to specify attack and defense actions at a high level (e.g., deploy a decoy), 2) an event-driven observability module that emits high-level observations (e.g., host is suspicious), 3) an environment state service to enable environment agnostic strategies, and 4) an attack graph service to reason how an attacker can explore an environment.
[…]
This reduction in effort can enable operators to rapidly 1) explore a variety of deception strategies against different attackers in different environments (see Sec. VII-A), 2) implement variations of strategies (see Sec. VII-B), and 3) evaluate extensions to consider new defense and attacker capabilities (see Sec. VII-B). Using Perry, we explore diverse scenarios and shed light on the efficacy of deceptive defenses (see Sec. VII-C). For instance,
@@
    We introduce a novel high-level and environment agnostic abstraction for expressing attackers and deception defenders: a state machine that executes high-level actions based on high-level observations of the environment (Sec. III).
@@
    We demonstrate the value of Perry by illustrating how it can enable operators to rapidly explore a wide variety of deception defenses, attackers, and environments, shedding light on the tradeoffs of deception defenses(Sec. VII).
Reproducible Research Perry is open-source and will be publicly available to the research community. [^1] The code is available at https://github.com/bsinger98/Perry.
@@ II-A The case for what-if experiments in deception
Following this example, if and how, could deception techniques have stopped the attack: Would some deception strategies have higher efficacy than others? To shed light on these questions, operators may want to explore what-if scenarios:
Example 1. The operator may want to quickly explore the efficacy of a wide variety of deception capabilities (e.g., decoy hosts [48], fake credentials [64], or fake software patches [6]) against many variants of this attacker.
Example 2. The operator may want to understand the benefits of adding better telemetry with the above deception algorithms, enabling the defense to react to attackers actions. For example, an operator may want to detect and react to attackers interacting with deceptive fake credentials the defender added.
@@ II-B Requirements and limitations of prior work
    Extensible: As new deception approaches, system capabilities, and attack strategies appear; the platform should be extensible to evaluate these new techniques.
At one extreme, we can consider simple agent-based simulators that are low effort [41, 30, 10, 61]. These do not emulate real networks, attackers, or defenders and lack the realism and expressiveness for real-world scenarios. We will later see in Sec. VII-A how subtle implementation details can greatly impact the efficacy of deception.
@@
Specifically, we consider a stateful defender (Example 2) that 1) deploys decoy hosts and credentials to the network, 2) detects hosts interacting with the decoy hosts and using the decoy credentials, and 3) restores any hosts that interact with decoy resources. Similarly, we create an attacker that mimics the real-world Equifax attacker [38] by: 1) gaining initial access to the network by infecting vulnerable web servers, 2) searching and using credentials to infect database hosts, and 3) exfiltrating data from each database host. Fig. 3 shows code snippets of the defender implementation.
Next, we highlight three issues with such implementations. [^2] Our goal is to illustrate problems broadly with available tooling, not at these specific tools. For instance, the problems we highlight with Caldera also exist in other attacker emulation tools [51, 25, 16] and the pain points of Elasticsearch with Python are similar to other defense tools [8, 27, 59].)
@@
    Low-level and cumbersome code: Existing tools offer low-level SDKs resulting in cumbersome low-level code. For instance, the stateful defender implementation requires 272 lines of low-level code. Deploying decoy hosts and credentials requires 213 lines of code using the Ansible and Openstack SDKs. The remaining 59 lines of code use the Elasticsearch API to detect attackers interacting with decoy hosts and credentials.
@@
    Non-portable code: Existing tools force the code to be tightly coupled with the environment, making it difficult to port attackers and defenders to a new environment. For instance, the attacker implementation exfiltrates data. In the Equifax environment, this requires (1) choosing a correct path of two hosts to exfiltrate the data on, and (2) choosing the correct exfiltration protocol between the hosts (e.g., the web server communicates with the database instance with SSH while the attacker and web server communicate with HTTPS). If the operator’s environment has a different path or different protocols, the attacker implementation will fail.
@@
    Modifications require distributed changes: As a consequence of complex low-level code, existing tools have complex dependencies between strategies, capabilities, or telemetry, making it difficult to extend to evaluate new deception approaches and attackers. Our implementations became tightly coupled to their capabilities and telemetry, making them difficult to extend. For instance, we may want to extend our deception to add decoy files to decoy hosts. But adding decoy files requires many distributed changes to the implementation. Specifically, in Fig. 3, we have to: (1) extend the initialization function to maintain lists of real and decoy file, (2) extend the detection rules to detect if fake data is being exfiltrated, (3) and (4) implement Ansible Playbooks to deploy the decoy files.
@@ III Perry System Overview
In this section, we give an overview of Perry. Perry introduces a high-level abstraction layer for security operators to specify attackers and deception defenders as well as an experimentation module to emulate these attackers and defenders in realistic network topologies.
@@
The inputs to Perry are the environment, attacker, and defender, shown in Fig. 4. Perry first instantiates the environment on a virtualized network and executes the defender and attacker logic expressed using its high-level abstraction. Perry collects metrics such as defender resources used (e.g., number of decoys used) and attacker success metrics (e.g., data exfiltrated) to enable quantitative analytics (Sec. VII).
@@ III-A High-level idea
To put the key ideas of Perry in context, let us revisit from first principles what the operator was trying to do and why existing tools failed to allow for the expression of those intents.
Conceptually, the operator has a mental model of the attacker and defender logic (Sec. II) shown in the simple state machines in Fig. 5. Here, the defender wants to deploy decoys, wait for attackers to interact with the decoys, and then runs actions to stop the attacker. Similarly, the attacker runs like a logical state machine that tries to infect web servers, find credentials, and finally exfiltrate data.
While these intents are easy to describe in this mental model, there is a significant disconnect with existing tools. Existing frameworks are environment-specific, low-level, and lack any abstraction, making it painful to realize these intents. For example, the first state of the defender in Fig. 5 is to deploy decoy hosts and credentials. But modeling this state in our implementation requires 240 lines of code (163 lines of Ansible code and 77 lines of Python code), which only work in one environment. As we discussed earlier, this code is tightly coupled, making it difficult to extend to new environments, capabilities, and strategies.
@@
Based on this insight, Perry introduces a high-level programming model, explicitly representing attackers and defenders as logical state machines that execute high-level actions based on high-level observations of the environment. In essence, Perry helps the user express high-level intentions of what they need to do, decoupling intentions from the how in low-level tools.
@@ III-B Anatomy of an experiment
Deception defense: In Perry, deception defense has three stages: A) initialization, set up strategies and deploy deception resources; B) event listeners, strategies subscribe to high-level observations from the telemetry service; and C) event handlers, strategies react to the high-level observations. For instance, the pseudo code of the stateful deception approach in Fig. 5 is (full example in Sec. V):
@@
# (B) Event listeners, detect interactions e.g.,
@@ IV Detailed system design
Next, we describe the key modules’ design in Perry: environment state service, action translator, observability module, and attack graph service. Broadly speaking, strategies subscribe to high-level observations from the observability module and output high-level actions for translation by the action translator, shown in Fig. 6.
@@ IV-A Environment state service
As shown in Sec. II, strategies in existing tools are coupled to a single environment. Rather than strategies having low-level and environment-specific code, in Perry, code that reasons about the environment is delegated to the environment state service. Perry’s modules and strategies query the environment state service for environment information (e.g., check if a credential is fake). To answer these queries, the environment state service maintains a network information base [25] by subscribing to high-level events from the observability module (Sec. IV-C). For example, for attackers, the environment state service subscribes to a new host discovery event and adds hosts to the knowledge base as they are discovered.
As an illustration, the following snippet shows a simple randomized decoy host and credential deployment using the Perry API. The key idea here is that all environment-specific code is delegated to the environment state service:
@@
In the above example, the implementation is now environment-agnostic. Rather than the code specifying low-level individual IP addresses to deploy decoy credentials, it uses the environment state service to select the high-level intention of the host for deployment of the decoy credentials.
@@ IV-B Action translator
Currently, actions in attack and defense emulation tools [5, 25, 16, 52] are specified with low-level commands that have pre and post-conditions. For example, one command might be an nmap scan script with the post-condition to execute an SSH exploit if a known CVE is found.
@@
In Perry, the high-level actions are decoupled from their implementations. As a result, the action translator makes it easier to specify, modify, and reason about strategies because users no longer have to reason about low-level code. Furthermore, the decoupling of implementations makes the action translator extensible in terms of both adding new high-level actions (e.g., adding a stealthy scan action) and adding new implementations of actions (e.g., a better data exfiltration technique). For example, in a strategy, we can program an exfiltrate data action as:
@@
We also design a library of common high-level deception and attack actions that are portable across environments in Table I. For instance, the data exfiltration action (1) uses the attack graph service to identify the shortest exfiltration path and (2) uses the environment state service to choose the correct protocol (e.g., HTTP) for data exfiltration.